4. Suppose a hospital is given a grant of 1.5 million dollars from the federal government for improvements in care. The managers at the hospital have to determine which departments receive what share of the grant. The following table is a proposed budgetary allocation for each of the main 4 departments at the hospital.

Department Proposed share of Grant
surgery $1,020,000.00
Emergency $255000.00
Pediatrics $150,000.00
Pharmacy $75,000.00

A) What proportion of the funding will potentially go to the pediatrics department?

B) What proportion of funding will go to all departments other than surgery?
Answer #1

a)
proportion of the funding will potentially go to the pediatrics department = 150000/1500000 = 0.1000

b)
proportion of funding will go to all departments other than surgery = (1500000 - 1020000)/1500000 = 0.3200